find the equation of a line that is perpendicular to the line y=1/6x+4 and contains the point (-3,- 0)
Answer by mananth(16946)

1 y = 1/ 6 x + 4
Divide by 1
y = 1/ 6 x + 4
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = 1/6

The slope of a line perpendicular to the above line will be the negative reciprocal -6
m1*m2=-1
The slope of the required line will be -6

m= -6 ,point ( -3 , 0 )
Find b by plugging the values of m & the point in
y=mx+b
0 = 18.00 + b
b= -18
m= -6
The required equation is y = -6 x-18
m.ananth@hotmail.ca